OpenStudy (anonymous):

is the total surface are S of a cube a function of the edge length x of the cube?

OpenStudy (ahsome):

Is the question: What is the Total Surface Area of a cube when the length is X?

OpenStudy (anonymous):

no, its asking if it is even a function to begin with. then it says: if so, write the function rule s(x) and evaluate s(3). and if not, why?

OpenStudy (ahsome):

\[s(x)=6x^2\] Where S is the Surface area of the cube Is that it? If so, substitute 3 for x and see the TSA

OpenStudy (ahsome):

I mean S is Surface Area, X is side value

OpenStudy (anonymous):

that sounds about right! thanks!

OpenStudy (ahsome):

sub value 3 into equation: \[s(x)=6x^2\] \[s(3)=6*3^2\] \[s(3)=6*9\] \[s(3)=54\]

OpenStudy (ahsome):

No problem :) If you think someone helped, just press the Best Response on the right.